OpenStudy (sasogeek):

2cos(6y)+11cos(6y)sin(3y) i'm stuck somewhere in the middle... not because I can't solve it... my calculator won't respond... please complete it for me. cos(6y)=0 or 2+11sin(3y)=0 <=that's where i'm stuck

OpenStudy (sasogeek):

please solve in radians

OpenStudy (sasogeek):

radians... please...

OpenStudy (sasogeek):

my calculator won't respond so please do all the work :) thank you

OpenStudy (lalaly):

ok sorry cos inverse of 0 is pi/2

OpenStudy (sasogeek):

i have the solutions... i just need to know how to get there

OpenStudy (ash2326):

\[\cos^{-1} 0=6y\] or 6y=(2n+1)*pi/2 y=(2n+1)*pi/12

OpenStudy (ash2326):

either this or 2+11 sin 3y=0 sin 3y =-2/11 \[3y=\sin^{-1} (-2/11 )\] \[3y=-\sin^{-1}(2/11)\] \[y=-(\sin^{-1} (2/11))/3\]

OpenStudy (sasogeek):

ok so i just opened up the page with the step-by-step solution... here's a snapshot

OpenStudy (ash2326):

cos 6y=0 6y=(2n+1)pi/2 n= 0, +1 or -1 and so on y=2npi/12+pi/12 y=n*pi/6+pi/12 y=pi/12 +pi*n/6

OpenStudy (sasogeek):

thanks btw, i'll go for a calc from a friend later today :) until then, thats all for now

OpenStudy (ash2326):

saso geek in the answer ,the first solution misses some answers, you might need to check it again with your instructor:)